OpenStudy (anonymous):

caitlin had $402 in her bank account she withdrew $15 each week to pay for a swimming lesson she now has $237 A. write an equation that can be used to find the number of swimming lessons that she paid for. B. at the time she has $237 the cost of a lesson rose to $19 how many lesson can she pay for with her remaining $237?

OpenStudy (anonymous):

this is free right?

OpenStudy (anonymous):

A. 402-15x=237

OpenStudy (bahrom7893):

obviously it's free! We're all volunteers haha

OpenStudy (bahrom7893):

The answer's 11

OpenStudy (anonymous):

B. 237-19x=0

OpenStudy (bahrom7893):

http://openstudy.com/users/bahrom7893#/updates/4f39bc2be4b0fc0c1a0e4e8f That's the exact same question someone answered earlier.

OpenStudy (anonymous):

i got 11 too...is it 12 or 11

OpenStudy (bahrom7893):

I mean i answered, someone asked haha it's 11

OpenStudy (anonymous):

okay what about c

OpenStudy (anonymous):

there is no c?

OpenStudy (anonymous):

this is C. At the time she has $237, the cost of the lesson rose to $19. How many lessons can she pay for with her remaining $237 ?

OpenStudy (anonymous):

She has 237 dollars. x lessons cost 19x dollars. Find the value at which 237-19x=0

OpenStudy (anonymous):

She can afford \[12\frac{9}{19}\] lessons, but since she can only purchase a full lesson or no lesson at all, she can only afford 12 lessons.

OpenStudy (anonymous):

okay thanks..Happy Valentines Day...lol:)